OpenStudy (ledah):

Solve by substitution x + 3y = -5 2x - y = 18

OpenStudy (anonymous):

x=-5-3y 2x-y=18 substitute with x 2(-5-3y)-y=18 -10-6y-y=18 -7y=18+10 -7y=28 y=-4 . Now go to any of the 2 equation and sbs. y in it and get x

OpenStudy (anonymous):